# Kineo: Calibration-Free Metric Motion Capture From Sparse RGB Cameras

### v1

@@
Abstract
Markerless multiview motion capture is often constrained by the need for precise camera calibration, limiting accessibility for non-experts and in-the-wild captures. Existing calibration-free approaches mitigate this requirement but suffer from high computational cost and reduced reconstruction accuracy.
@@ Abstract
We present Kineo, a fully automatic, calibration-free pipeline for markerless motion capture from videos captured by unsynchronized, uncalibrated, consumer-grade RGB cameras. Kineo leverages 2D keypoints from off-the-shelf detectors to simultaneously calibrate cameras, including Brown–Conrady distortion coefficients, and reconstruct 3D keypoints and dense scene point maps at metric scale. A confidence-driven spatio-temporal keypoint sampling strategy, combined with graph-based global optimization, ensures robust calibration at a fixed computational cost independent of sequence length. We further introduce a pairwise reprojection consensus score to quantify 3D reconstruction reliability for downstream tasks.
Evaluations on EgoHumans and Human3.6M demonstrate substantial improvements over prior calibration-free methods. Compared to previous state-of-the-art approaches, Kineo reduces camera translation error by approximately 83–85%, camera angular error by 86–92%, and world mean-per-joint error (W-MPJPE) by 83–91%.
Kineo is also efficient in real-world scenarios, processing multi-view sequences faster than their duration in specific configuration (e.g., 36min to process 1h20min of footage). The full pipeline and evaluation code are openly released to promote reproducibility and practical adoption at: liris-xr.github.io/kineo.
1 Introduction
@@
Motion capture technology is an essential tool for converting physical movement into digital representations. Traditional optical systems (e.g., OptiTrack [30], Vicon [51], Qualisys [36]) rely on multiple calibrated infrared cameras that track reflective markers placed on key locations of the subject’s body. An animated skeletal model is then reconstructed for the motion to be analyzed or re-targeted on a virtual character for a realistic animation. While these systems achieve sub-millimeter accuracy in controlled studio conditions, they require subjects to wear optical markers, which can induce behavioral biases and make their use challenging or impractical. Moreover, deployment for in-the-wild scenarios (e.g., outdoor sport, theatrical performances or concerts, craftsmanship), or for casual users (e.g., home fitness, motion-based gaming) presents significant challenges: the setup requires careful calibration, controlled lighting with limited infrared perturbations, and precise marker placement on the body. All cameras are wired together for temporal synchronization and connected to power outlets, limiting their use for on-location recordings in uncontrolled environments. Furthermore, the specialized hardware and controlled studio conditions significantly restrict the accessibility of these systems due to their high cost.
@@
Calibration-free methods have been explored to address these challenges by recovering camera parameters post-capture using video and audio information, eliminating the need for calibration procedures. While great progress has been made recently, especially in reconstructing the cameras, humans, and scene in metric scale [29, 38], calibration-free approaches remain relatively underexplored, and there is still significant potential for improvement in terms of accuracy, scalability to long sequences by reducing computational cost, and practicality for downstream tasks. In particular, current methods suffer from pipeline inefficiencies such as heavy reliance on multiple computationally intensive neural network models, the use of parametric model representations and iterative nonlinear optimization on individual frames. These factors collectively limit the scalability to long video sequences on commodity hardware. Moreover, many approaches lack 3D confidence measures, which are essential for accurately quantifying uncertainty in the reconstructed geometry.
@@
  • Confidence‑Aware Spatio-Temporal Subsampling. We introduce a novel keypoint correspondence spatio-temporal sampling strategy that uses 2D detector confidences to select high‑quality matches for robust and more efficient RANSAC‑based Structure-from-Motion for essential‑matrix estimation.
@@
  • Model-Agnostic and SMPL-Based Scale Recovery. To address global scale ambiguity, we evaluate two independent approaches: (1) a model-agnostic method using monocular metric depth estimation, and (2) a human-prior-based method that leverages the SMPL model, similar to HSfM, when people are present in the scene.
@@ 2.1 Temporal Synchronization
Temporal synchronization refers to the process of recovering the time offset between the camera internal clock and a global time reference to express the timestamps of each frame in a shared world time reference. Temporal synchronization is essential in multi-view motion capture to synchronize camera frames, as even small temporal offsets between cameras can lead to inaccuracies in the reconstruction due to visual differences in the observed poses across views, especially when dealing with fast movements [44, 5, 52]. Traditional studio camera systems rely on physical wiring to trigger all cameras simultaneously. However, such hardware-based synchronization is not available or impractical when using consumer-grade devices like smartphones or action cameras. Alternatively, jam syncing can be used. This technique synchronizes the internal clocks of multiple recording devices by transferring a timecode from a master device to slave devices. The synchronization must be performed at the start of the recording and periodically throughout the session to prevent clock drift. Once the slave devices receive the timecode, they operate independently while remaining aligned with the master device’s time reference. Importantly, jam syncing requires dedicated hardware [46] or software [24] at capture time. When dedicated synchronization hardware or software is unavailable or was not used during capture, post-capture temporal synchronization methods can be employed. Among these, audio-based techniques have been used [43, 11]. These methods leverage the sound captured by each camera to find the time offsets (temporal extrinsics). Hasler et al. [11] use cross-correlation on the audio signal using Fast Fourier Transform; the audio delay between the signals is found by locating the peak of the cross-correlation. When cameras are located at varying distances from the sound source, differences in audio propagation time can cause desynchronization. For instance, a 20-meter distance difference between two cameras recording at 60 Hz would lead to an estimated misalignment of around 4 frames. However, such large disparities are rare in practice, since the sound is typically emitted by the subject, who is usually positioned near the center of a circular or spherical camera setup. In these configurations, the propagation delay is negligible. Even in this worst-case scenario, the resulting 4-frame misalignment can still serve as a coarse initial estimate for image-based synchronization methods. These methods use visual information, typically human pose, from the different cameras to refine the time offset but assume that an initial coarse temporal alignment is provided for the optimization. Shuai et al. [44] include the time offset as a variable to be optimized within a joint estimation framework, assuming identical frame rates between all cameras. Vo et al. [52] optimize frame rate, rolling shutter pixel readout speed and time offset. To perform sub-frame alignment, those methods interpolate poses to a common time frame for their optimization process. Shuai et al. [44] estimate joint velocities using finite differences, allowing them to interpolate poses using a first-order approximation. While computationally efficient, this approach is sensitive to noise. To improve robustness, Akhter et al. [1] proposed using Discrete Cosine Transform (DCT) for interpolating the values, which is less sensitive to noise. This strategy was subsequently adopted in later works [52, 15]. More recently, Choi et al. [6] proposed using dynamic time warping (DTW) between every pair of videos to compute a constant time offset that minimizes the distance between 3D joint positions, extracted from estimated SMPL models obtained via SLAHMR [61] for each video individually. Their method achieves subframe accuracy without requiring coarse alignment. However, it involves constructing an $N\times N$ matrix containing all possible pose associations to find the optimal time offset that minimizes the 3D joint distances, which can become a computational bottleneck for large sequences.
@@
2.2 Automatic Camera Calibration
Accurate estimation of camera parameters, both intrinsics (e.g., focal length, principal point, distortion coefficients) and extrinsics (position and orientation), is essential for multi-view 3D reconstruction. Examples include the use of such parameters in methods that first regress 2D keypoints or features and subsequently lift them into 3D space through techniques such as triangulation [5, 62, 56] or volumetric fusion [17, 49, 60, 35]. In optimization-based pipelines, the estimated camera parameters are used to reproject 3D structures (commonly a SMPL model or skeletal representation) into 2D space, minimizing the reprojection error with respect to the observed 2D cues. While these methods exploit accurate camera parameters, most assume access to ground-truth values provided by datasets or obtained through manual calibration, rather than estimating them. This section focuses on existing methods for automatically estimating camera intrinsics and extrinsics.
2.2.1 Camera Intrinsics
@@
High-fidelity intrinsic calibration is essential in both monocular and multi-view systems, as it directly impacts the accuracy of the 2D–3D mapping through the projection operation. In monocular regression-based approaches that predict 3D structure directly from 2D inputs, having access to precise intrinsics at inference time has shown to improve the quality of the reconstruction [8, 20, 32, 39, 3]. Incorporating intrinsics into the network inputs enables more generalizable performance compared to methods using a simplified model of orthographic projection or fixed training-set intrinsics. These simplified models introduce a camera bias, resulting in a trade-off: achieving more accurate 3D pose estimates can come at the expense of poorer alignment with 2D image features, and vice versa [8, 20].
@@
In multi-view reconstruction, intrinsics are equally critical. Both in methods where intrinsics are used to unproject the 2D keypoints or features into a shared 3D space [17, 49, 60, 35] or in optimization-based methods when computing the reprojection loss to align the 3D model with 2D cues [13, 29].
@@
Because camera intrinsics are fixed properties of the camera itself and independent of its viewpoint, they can be obtained through offline calibration after data capture. Consequently, some methods relax the constraints on obtaining these parameters by using ground truth intrinsics, assuming the camera is accessible for later calibration and thus only focus on recovering the camera extrinsics [56, 13, 6]. However, this assumption limits applicability in scenarios involving online videos, unknown camera setups, or fully calibration-free approaches.
@@
Several learning-based methods have been proposed to enable calibration-free inference by estimating camera parameters directly from image data. Within the Human Pose Estimation domain, some monocular approaches focus on estimating the camera field of view from images containing humans [32, 20] to feed them to the body shape and pose regression model or for later optimization.
@@
More recently, multiview approaches have also incorporated intrinsic parameter estimation. For instance, HAMSt3R [38] builds upon DUSt3R [55] and MASt3R [25], recovering intrinsics from the predicted depth maps but with a model specialized for human-centric reconstruction which yields better reconstruction quality. Similarly, Müller et al. [29] leverage MASt3R to initialize camera intrinsics that are subsequently refined in a global optimization loop. Both methods remain tailored to human subjects and do not estimate distortions coefficients. More broadly, other learning-based methods have been introduced for general-purpose applications, predicting both the focal length and lens distortion parameters [50, 47, 4], or just the camera focal length [54, 25, 55, 54] from RGB images. In our method, we adopt the general model MoGe [54] to predict an initial camera focal length for each view (see Sec. 5.1). The focal length is subsequently optimized in the bundle adjustment step (see Sec. 5.6). To our knowledge, Human Pose Estimation methods, whether monocular or multiview, typically assume distortion-free inputs, relying on ground-truth calibration to undistort images as a preprocessing step prior to reconstruction or evaluation or simply neglecting distortion altogether. In contrast, our pipeline also estimates distortion coefficients, following the Brown–Conrady model, by integrating them in the bundle adjustment (see Sec. 5.6), enabling accurate handling of lens distortion and improving multi-view consistency.
2.2.2 Camera Extrinsics
@@
Learning-based human-centric methods often leverage strong priors on human shape and pose. Choi et al. [6] first estimate monocular human motion in each view using SLAHMR [61], which predicts SMPL body meshes per frame. They compute an initial extrinsic estimate by minimizing per-joint distances to align motions across views, followed by refinement using a NeRF-based photometric optimization. This achieves sub-degree rotational and sub-centimeter translational errors on the CMU Panoptic Studio dataset.
@@
Hewitt et al. [14] adopt a PnP-based strategy, aligning 3D head meshes with 2D facial landmarks, following the method introduced by Szymanowicz et al. [45]. The head serves as a stable reference as its overall shape remains relatively stable across facial expressions and head poses. Initial estimates are refined through direct optimization of camera rotations and translations.
@@
Müller et al. [29] use DUSt3R [55] to predict dense 3D pointmaps from image pairs, recover relative poses, and integrate them into a global optimization that simultaneously updates SMPL models, scene pointmaps, and camera parameters. Similarly, HAMSt3R [38], trained specifically on human-centric scenes, aligns multiple views using predicted pairwise pointmaps within a common global frame, based on the method from DUSt3R [55].
@@
Despite differences, these methods share a reliance on parametric body models (e.g., SMPL or SOMA) or human-specific scene training. Such priors provide strong constraints that help resolve global scale ambiguities, as body shape enforces a consistent metric scale, but are computationally intensive, especially when optimization to refine initial camera poses and body poses is involved, like in HSfM [29].
@@
Classical SfM approaches estimate extrinsics by exploiting geometric correspondences between views. Frameworks like COLMAP [40] perform incremental bundle adjustment using traditional feature matching. While computationally efficient, these methods are sensitive to noise, outliers, and unreliable correspondences.
@@
Outlier rejection is commonly implemented by embedding the estimation algorithm within a RANSAC framework [34, 56], which robustly estimates relative camera poses from noisy keypoint correspondences. However, uniform random sampling in long sequences is often inefficient: matches from unreliable regions may be repeatedly sampled, while spatially clustered keypoints contribute redundant geometric information. To address this, we introduce a confidence-weighted spatio-temporal subsampling strategy that prioritizes high-confidence, spatially diverse keypoints, ensuring robustness and fixed computational cost.
@@
Early motion-capture SfM approaches also exploited semantic cues from the human body, including silhouette extrema [48, 28, 22], bounding box centers [57], and 2D keypoints [34, 56], with keypoints proving particularly robust for establishing correspondences.
@@
UnCaliPose [56] follows a similar principle, using 2D keypoints to estimate initial camera poses that are later refined through joint bundle adjustment with 3D human poses. Although effective, the method depends on ground-truth intrinsics and distortion coefficients, does not reconstruct the scene, and shows limited accuracy in metric-scale outputs. To improve calibration, UnCaliPose adopts a two-step approach, first calibrating cameras across multiple frames and then solving for human poses, but the paper does not specify how these frames are selected. Without a selection strategy, incorporating all frames makes joint bundle adjustment computationally infeasible for long sequences. Our approach builds upon this framework by also leveraging 2D keypoints, but extends it to estimate camera intrinsics and distortion parameters, reconstruct a scene pointmap (as in HSfM and HAMSt3R), and produce metric-scale results with confidence estimates. Our spatio-temporal keypoint sampling ensures both high quality automatic calibration and scalable computational cost during bundle adjustment.
@@
Once pairwise relative poses are estimated using SfM, the next step is to recover the global camera calibration by composing these pairwise transformations into a consistent network. Greedy strategies, such as selecting the next camera based on maximum inliers [34] or minimal average alignment angle [56], offer efficiency, but accumulate errors and provide a suboptimal solution. In addition, relative extrinsics computed via SfM are only determined up to an unknown scale factor, which must be resolved to properly compose them into the absolute extrinsics. Xu et al. [56] address this by leveraging the fixed length of bones in 3D space to resolve scale ambiguities between camera pairs. However, this approach requires that the bone is visible and well triangulated by all camera pairs involved, which may not always be the case in challenging capture scenarios.
@@
We reformulate calibration as a graph problem: cameras are nodes, pairwise calibrations forming the edges are weighted by a pairwise calibration score, and a minimum spanning tree yields the globally optimal calibration. This formulation, that can be found in literature as the essential graph, allows us to solve for the relative scales via loop closure constraints, and to find the optimal camera network configuration via the use of a minimum spanning tree. To recover the global scale, we explore two approaches. The first leverages a monocular metric geometry estimator, MoGe [54] that we use to predict per-view depth maps in real-world units. This allows the reconstruction to be rescaled accurately and removes the need for any category-specific parametric model. The second approach draws inspiration from prior work using human body models; we estimate global scale as a downstream step using the shape priors of the SMPL model. Both strategies preserve the efficiency of a classical Structure-from-Motion pipeline while enabling metric-scale reconstruction. As a result, our method produces calibrated cameras, 3D keypoints, and a point cloud of the surrounding scene, matching the outputs of previous works.
2.3 3D Confidence
@@
To overcome this, probabilistic approaches have been introduced to model uncertainty by predicting distributions over plausible 3D poses. For example, ProHMR [21] and HumaniFlow [42] sample multiple hypotheses from learned distributions, allowing per-joint or per-vertex uncertainty to be estimated from the variance of the samples. These uncertainty estimates enable view-consistent fusion—either by optimizing the joint probability across views or by weighting individual predictions in a fusion step according to their confidence. However, this expressiveness comes at the cost of computational efficiency, as accurate uncertainty estimation typically requires drawing a large number of samples.
More efficient alternatives such as POCO [7] have been proposed to address this trade-off. POCO extends deterministic models by directly predicting uncertainty alongside the 3D output, eliminating the need for sampling. While significantly faster, such methods remain tightly coupled to parametric models and are not readily applicable to non-parametric 3D representations such as lifted keypoints.

### v2

@@
Abstract.
@@
Markerless multiview motion capture is often constrained by the need for precise camera calibration, limiting accessibility for non-experts and in-the-wild captures. Existing calibration-free approaches mitigate this requirement but suffer from high computational cost and reduced reconstruction accuracy. We present Kineo, a fully automatic, calibration-free pipeline for markerless motion capture from videos captured by unsynchronized, uncalibrated, consumer-grade RGB cameras. Kineo leverages 2D keypoints from off-the-shelf detectors to simultaneously calibrate cameras, including Brown-Conrady distortion coefficients, and reconstruct 3D keypoints and dense scene point maps at metric scale. A confidence-driven spatio-temporal keypoint sampling strategy, combined with graph-based global optimization, ensures robust calibration at a fixed computational cost independent of sequence length. We further introduce a pairwise reprojection consensus score to quantify 3D reconstruction reliability for downstream tasks. Evaluations on EgoHumans and Human3.6M demonstrate substantial improvements over prior calibration-free methods. Compared to previous state-of-the-art approaches, Kineo reduces camera translation error by approximately 83-85%, camera angular error by 86-92%, and world mean-per-joint error (W-MPJPE) by 83-91%. Kineo is also efficient in real-world scenarios, processing multi-view sequences faster than their duration in specific configuration (e.g., 36min to process 1h20min of footage). The full pipeline and evaluation code are openly released to promote reproducibility and practical adoption at: liris-xr.github.io/kineo.
human pose estimation, motion capture, markerless, calibration-free
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Computing methodologies Reconstruction
1. Introduction
@@
Motion capture technology is an essential tool for converting physical movement into digital representations. Traditional optical systems (e.g., OptiTrack (NaturalPoint, 2025), Vicon (Vicon Motion Systems, 2025), Qualisys (Qualisys, 2025)) rely on multiple calibrated infrared cameras that track reflective markers placed on key locations of the subject’s body. An animated skeletal model is then reconstructed for the motion to be analyzed or re-targeted on a virtual character for a realistic animation. While these systems achieve sub-millimeter accuracy in controlled studio conditions, they require subjects to wear optical markers, which can induce behavioral biases and make their use challenging or impractical. Moreover, deployment for in-the-wild scenarios (e.g., outdoor sport, theatrical performances or concerts, craftsmanship), or for casual users (e.g., home fitness, motion-based gaming) presents significant challenges: the setup requires careful calibration, controlled lighting with limited infrared perturbations, and precise marker placement on the body. All cameras are wired together for temporal synchronization and connected to power outlets, limiting their use for on-location recordings in uncontrolled environments. Furthermore, the specialized hardware and controlled studio conditions significantly restrict the accessibility of these systems due to their high cost.
@@
Calibration-free methods have been explored to address these challenges by recovering camera parameters post-capture using video and audio information, eliminating the need for calibration procedures. While great progress has been made recently, especially in reconstructing the cameras, humans, and scene in metric scale (Müller et al., 2025; Rojas et al., 2025), calibration-free approaches remain relatively underexplored, and there is still significant potential for improvement in terms of accuracy, scalability to long sequences by reducing computational cost, and practicality for downstream tasks. In particular, current methods suffer from pipeline inefficiencies such as heavy reliance on multiple computationally intensive neural network models, the use of parametric model representations and iterative nonlinear optimization on individual frames. These factors collectively limit the scalability to long video sequences on commodity hardware. Moreover, many approaches lack 3D confidence measures, which are essential for accurately quantifying uncertainty in the reconstructed geometry.
@@
  • Confidence‑Aware Spatio-Temporal Subsampling. We introduce a novel keypoint correspondence spatio-temporal sampling strategy that uses 2D detector confidences to select high‑quality matches for robust and more efficient RANSAC based Structure-from-Motion for essential‑matrix estimation.
@@
  • Model-Agnostic and SMPL-Based Scale Recovery. To address global scale ambiguity, we evaluate two independent approaches: (1) a model-agnostic method using monocular metric depth estimation, and (2) a human-prior-based method that leverages the SMPL model when people are present in the scene.
@@ 2.1. Temporal Synchronization
Temporal synchronization refers to the process of recovering the time offset between the camera internal clock and a global time reference to express the timestamps of each frame in a shared world time reference. Temporal synchronization is essential in multi-view motion capture to synchronize camera frames, as even small temporal offsets between cameras can lead to inaccuracies in the reconstruction due to visual differences in the observed poses across views, especially when dealing with fast movements (Shuai et al., 2022; Bragagnolo et al., 2025; Vo et al., 2022). Traditional studio camera systems rely on physical wiring to trigger all cameras simultaneously. However, such hardware-based synchronization is not available or impractical when using consumer-grade devices like smartphones or action cameras. Alternatively, jam syncing can be used. This technique synchronizes the internal clocks of multiple recording devices by transferring a timecode from a master device to slave devices. The synchronization must be performed at the start of the recording and periodically throughout the session to prevent clock drift. Once the slave devices receive the timecode, they operate independently while remaining aligned with the master device’s time reference. Importantly, jam syncing requires dedicated hardware (Tentacle Sync, 2025) or software (Latimer et al., 2015) at capture time. When dedicated synchronization hardware or software is unavailable or was not used during capture, post-capture temporal synchronization methods can be employed. Among these, audio-based techniques have been used (Shrstha et al., 2007; Hasler et al., 2009). These methods leverage the sound captured by each camera to find the time offsets (temporal extrinsics). Hasler et al. (Hasler et al., 2009) use cross-correlation on the audio signal using Fast Fourier Transform; the audio delay between the signals is found by locating the peak of the cross-correlation. When cameras are located at varying distances from the sound source, differences in audio propagation time can cause desynchronization. For instance, a 20-meter distance difference between two cameras recording at 60 Hz would lead to an estimated misalignment of around 4 frames. However, such large disparities are rare in practice, since the sound is typically emitted by the subject, who is usually positioned near the center of a circular or spherical camera setup. In these configurations, the propagation delay is negligible. Even in this worst-case scenario, the resulting 4-frame misalignment can still serve as a coarse initial estimate for image-based synchronization methods. These methods use visual information, typically human pose, from the different cameras to refine the time offset but assume that an initial coarse temporal alignment is provided for the optimization. Shuai et al. (Shuai et al., 2022) include the time offset as a variable to be optimized within a joint estimation framework, assuming identical frame rates between all cameras. Vo et al. (Vo et al., 2022) optimize frame rate, rolling shutter pixel readout speed and time offset. To perform sub-frame alignment, those methods interpolate poses to a common time frame for their optimization process. Shuai et al. (Shuai et al., 2022) estimate joint velocities using finite differences, allowing them to interpolate poses using a first-order approximation. While computationally efficient, this approach is sensitive to noise. To improve robustness, Akhter et al. (Akhter et al., 2012) proposed using Discrete Cosine Transform (DCT) for interpolating the values, which is less sensitive to noise. This strategy was subsequently adopted in later works (Vo et al., 2022; Huang et al., 2017). More recently, Choi et al. (Choi et al., 2025) proposed using dynamic time warping (DTW) between every pair of videos to compute a constant time offset that minimizes the distance between 3D joint positions, extracted from estimated SMPL models obtained via SLAHMR (Ye et al., 2023) for each video individually. Their method achieves subframe accuracy without requiring coarse alignment. However, it involves constructing an $N\times N$ matrix containing all possible pose associations to find the optimal time offset that minimizes the 3D joint distances, which can become a computational bottleneck for large sequences.
@@
2.2. Automatic Camera Calibration
Accurate estimation of camera parameters, both intrinsics (e.g., focal length, principal point, distortion coefficients) and extrinsics (position and orientation), is essential for multi-view 3D reconstruction. Examples include the use of such parameters in methods that first regress 2D keypoints or features and subsequently lift them into 3D space through techniques such as triangulation (Bragagnolo et al., 2025; Zhang et al., 2021; Xu and Kitani, 2022) or volumetric fusion (Iskakov et al., 2019; Tu et al., 2020; Ye et al., 2022; Qiu et al., 2019). In optimization-based pipelines, the estimated camera parameters are used to reproject 3D structures (commonly a SMPL model or skeletal representation) into 2D space, minimizing the reprojection error with respect to the observed 2D cues. While these methods exploit accurate camera parameters, most assume access to ground-truth values provided by datasets or obtained through manual calibration, rather than estimating them. This section focuses on existing methods for automatically estimating camera intrinsics and extrinsics.
2.2.1. Camera Intrinsics
@@
High-fidelity intrinsic calibration is essential in both monocular and multi-view systems, as it directly impacts the accuracy of the 2D–3D mapping through the projection operation. In monocular regression-based approaches that predict 3D structure directly from 2D inputs, having access to precise intrinsics at inference time has shown to improve the quality of the reconstruction (Dwivedi et al., 2024b; Kocabas et al., 2021; Patel and Black, 2025; Sárándi and Pons-Moll, 2025; Baradel et al., 2025). Incorporating intrinsics into the network inputs enables more generalizable performance compared to methods using a simplified model of orthographic projection or fixed training-set intrinsics. These simplified models introduce a camera bias, resulting in a trade-off: achieving more accurate 3D pose estimates can come at the expense of poorer alignment with 2D image features, and vice versa (Dwivedi et al., 2024b; Kocabas et al., 2021).
@@
In multi-view reconstruction, intrinsics are equally critical. Both in methods where intrinsics are used to unproject the 2D keypoints or features into a shared 3D space (Iskakov et al., 2019; Tu et al., 2020; Ye et al., 2022; Qiu et al., 2019) or in optimization-based methods when computing the reprojection loss to align the 3D model with 2D cues (Hewitt et al., 2024a; Müller et al., 2025).
@@
Because camera intrinsics are fixed properties of the camera itself and independent of its viewpoint, they can be obtained through offline calibration after data capture. Consequently, some methods relax the constraints on obtaining these parameters by using ground truth intrinsics, assuming the camera is accessible for later calibration and thus only focus on recovering the camera extrinsics (Xu and Kitani, 2022; Hewitt et al., 2024a; Choi et al., 2025). However, this assumption limits applicability in scenarios involving online videos, unknown camera setups, or fully calibration-free approaches.
@@
Several learning-based methods have been proposed to enable calibration-free inference by estimating camera parameters directly from image data. Within the Human Pose Estimation domain, some monocular approaches focus on estimating the camera field of view from images containing humans (Patel and Black, 2025; Kocabas et al., 2021) to feed them to the body shape and pose regression model or for later optimization.
@@
More recently, multiview approaches have also incorporated intrinsic parameter estimation. For instance, HAMSt3R (Rojas et al., 2025) builds upon DUSt3R (Wang et al., 2024) and MASt3R (Leroy et al., 2024), recovering intrinsics from the predicted depth maps but with a model specialized for human-centric reconstruction which yields better reconstruction quality. Similarly, Müller et al. (Müller et al., 2025) leverage MASt3R to initialize camera intrinsics that are subsequently refined in a global optimization loop. Both methods remain tailored to human subjects and do not estimate distortions coefficients. More broadly, other learning-based methods have been introduced for general-purpose applications, predicting both the focal length and lens distortion parameters (Veicht et al., 2024; Tirado-Garín and Civera, 2025; Bogdan et al., 2018), or just the camera focal length  (Wang et al., 2025b; Leroy et al., 2024; Wang et al., 2024, 2025b) from RGB images. In our method, we adopt the general model MoGe (Wang et al., 2025b) to predict an initial camera focal length for each view (see Sec. 5.1). The focal length is subsequently optimized in the bundle adjustment step (see Sec. 5.6). To our knowledge, Human Pose Estimation methods, whether monocular or multiview, typically assume distortion-free inputs, relying on ground-truth calibration to undistort images as a preprocessing step prior to reconstruction or evaluation or simply neglecting distortion altogether. In contrast, our pipeline also estimates distortion coefficients, following the Brown–Conrady model, by integrating them in the bundle adjustment (see Sec. 5.6), enabling accurate handling of lens distortion and improving multi-view consistency.
2.2.2. Camera Extrinsics
@@
Learning-based human-centric methods often leverage strong priors on human shape and pose. Choi et al. (Choi et al., 2025) first estimate monocular human motion in each view using SLAHMR (Ye et al., 2023), which predicts SMPL body meshes per frame. They compute an initial extrinsic estimate by minimizing per-joint distances to align motions across views, followed by refinement using a NeRF-based photometric optimization. This achieves sub-degree rotational and sub-centimeter translational errors on the CMU Panoptic Studio dataset.
@@
Hewitt et al. (Hewitt et al., 2024b) adopt a PnP-based strategy, aligning 3D head meshes with 2D facial landmarks, following the method introduced by Szymanowicz et al. (Szymanowicz et al., 2022). The head serves as a stable reference as its overall shape remains relatively stable across facial expressions and head poses. Initial estimates are refined through direct optimization of camera rotations and translations.
@@
Müller et al. (Müller et al., 2025) use DUSt3R (Wang et al., 2024) to predict dense 3D pointmaps from image pairs, recover relative poses, and integrate them into a global optimization that simultaneously updates SMPL models, scene pointmaps, and camera parameters. Similarly, HAMSt3R (Rojas et al., 2025), trained specifically on human-centric scenes, aligns multiple views using predicted pairwise pointmaps within a common global frame, based on the method from DUSt3R (Wang et al., 2024).
@@
Despite differences, these methods share a reliance on parametric body models (e.g., SMPL or SOMA) or human-specific scene training. Such priors provide strong constraints that help resolve global scale ambiguities, as body shape enforces a consistent metric scale, but are computationally intensive, especially when optimization to refine initial camera poses and body poses is involved, like in HSfM (Müller et al., 2025).
@@
Classical SfM approaches estimate extrinsics by exploiting geometric correspondences between views. Frameworks like COLMAP (Schönberger and Frahm, 2016) perform incremental bundle adjustment using traditional feature matching. While computationally efficient, these methods are sensitive to noise, outliers, and unreliable correspondences.
@@
Outlier rejection is commonly implemented by embedding the estimation algorithm within a RANSAC framework (Puwein et al., 2015; Xu and Kitani, 2022), which robustly estimates relative camera poses from noisy keypoint correspondences. However, uniform random sampling in long sequences is often inefficient: matches from unreliable regions may be repeatedly sampled, while spatially clustered keypoints contribute redundant geometric information. To address this, we introduce a confidence-weighted spatio-temporal subsampling strategy that prioritizes high-confidence, spatially diverse keypoints, ensuring robustness and fixed computational cost.
@@
Early motion-capture SfM approaches also exploited semantic cues from the human body, including silhouette extrema (Tsuhan Chen et al., 2007; Lv et al., 2006; Krahnstoever and Mendonca, 2005), bounding box centers (Xu et al., 2021), and 2D keypoints (Puwein et al., 2015; Xu and Kitani, 2022), with keypoints proving particularly robust for establishing correspondences.
@@
UnCaliPose (Xu and Kitani, 2022) follows a similar principle, using 2D keypoints to estimate initial camera poses that are later refined through joint bundle adjustment with 3D human poses. Although effective, the method depends on ground-truth intrinsics and distortion coefficients, does not reconstruct the scene, and shows limited accuracy in metric-scale outputs. To improve calibration, UnCaliPose adopts a two-step approach, first calibrating cameras across multiple frames and then solving for human poses, but the paper does not specify how these frames are selected. Without a selection strategy, incorporating all frames makes joint bundle adjustment computationally infeasible for long sequences. Our approach builds upon this framework by also leveraging 2D keypoints, but extends it to estimate camera intrinsics and distortion parameters, reconstruct a scene pointmap (as in HSfM and HAMSt3R), and produce metric-scale results with confidence estimates. Our spatio-temporal keypoint sampling ensures both high quality automatic calibration and scalable computational cost during bundle adjustment.
@@
Once pairwise relative poses are estimated using SfM, the next step is to recover the global camera calibration by composing these pairwise transformations into a consistent network. Greedy strategies, such as selecting the next camera based on maximum inliers (Puwein et al., 2015) or minimal average alignment angle (Xu and Kitani, 2022), offer efficiency, but accumulate errors and provide a suboptimal solution. In addition, relative extrinsics computed via SfM are only determined up to an unknown scale factor, which must be resolved to properly compose them into the absolute extrinsics. Xu et al. (Xu and Kitani, 2022) address this by leveraging the fixed length of bones in 3D space to resolve scale ambiguities between camera pairs. However, this approach requires that the bone is visible and well triangulated by all camera pairs involved, which may not always be the case in challenging capture scenarios.
@@
We reformulate calibration as a graph problem: cameras are nodes, pairwise calibrations forming the edges are weighted by a pairwise calibration score, and a minimum spanning tree yields the globally optimal calibration. This formulation, that can be found in literature as the essential graph, allows us to solve for the relative scales via loop closure constraints, and to find the optimal camera network configuration via the use of a minimum spanning tree. To recover the global scale, we explore two approaches. The first leverages a monocular metric geometry estimator, MoGe (Wang et al., 2025b) that we use to predict per-view depth maps in real-world units. This allows the reconstruction to be rescaled accurately and removes the need for any category-specific parametric model. The second approach draws inspiration from prior work using human body models; we estimate global scale as a downstream step using the shape priors of the SMPL model. Both strategies preserve the efficiency of a classical Structure-from-Motion pipeline while enabling metric-scale reconstruction. As a result, our method produces calibrated cameras, 3D keypoints, and a point cloud of the surrounding scene, matching the outputs of previous works.
2.3. 3D Confidence
@@
To overcome this, probabilistic approaches have been introduced to model uncertainty by predicting distributions over plausible 3D poses. For example, ProHMR (Kolotouros et al., 2021) and HumaniFlow (Sengupta et al., 2023) sample multiple hypotheses from learned distributions, allowing per-joint or per-vertex uncertainty to be estimated from the variance of the samples. These uncertainty estimates enable view-consistent fusion—either by optimizing the joint probability across views or by weighting individual predictions in a fusion step according to their confidence. However, this expressiveness comes at the cost of computational efficiency, as accurate uncertainty estimation typically requires drawing a large number of samples.
More efficient alternatives such as POCO (Dwivedi et al., 2024a) have been proposed to address this trade-off. POCO extends deterministic models by directly predicting uncertainty alongside the 3D output, eliminating the need for sampling. While significantly faster, such methods remain tightly coupled to parametric models and are not readily applicable to non-parametric 3D representations such as lifted keypoints.